OpenStudy (anonymous):

Solve for x: 2.05 = x / .16-x

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

im having trouble could you help me??

OpenStudy (anonymous):

it's a 2 plus .05 or 2 point .05

OpenStudy (anonymous):

its a decimal 2.05

OpenStudy (anonymous):

(0.16-x)2.05=x get it??

OpenStudy (anonymous):

then 0.328-0.05x=x

OpenStudy (anonymous):

how did you do that miamy??

OpenStudy (anonymous):

i get 0.328 - 2.05x = x

OpenStudy (anonymous):

so which part u did not get it yet

OpenStudy (anonymous):

im stuck on 0.328 - 2.05x = x

OpenStudy (anonymous):

ok.........then -2.05x=x-0.328

OpenStudy (anonymous):

alright i get that but then what do you do?

OpenStudy (anonymous):

nonnooo.......easier way......sorry.....then 0.328=x+2.05x

OpenStudy (anonymous):

0.328=3.05x

OpenStudy (anonymous):

how did you do that??

OpenStudy (anonymous):

oh wait i understand

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

i got it thank you its very easy now

OpenStudy (anonymous):

x=0.328/3.05

OpenStudy (anonymous):

alright..uw